How much can Haydon Rheo systems(F) , Inc., afford to spend now on an energy management system if the software will save the company $21,300(A) per year for the next 5 years? Use an interest rate of 10% per year
Answer #1

Annual savings = 21,300

Time period = 5 years

Interest Rate = 10%

How much the company afford to spend now?
For this calculate the PRESENT VALUE of the annual savings.
We have to use the (P/A, I, N) factor to calculate the present value of annual savings.

Present Value = A(P/A, 10%, 5)

Present Value = 21,300 (P/A, 10%,5)

Present Value = 21,300 (3.79079)=80,743.8

Rounding to nearest dollar PW = $80,744

The company afford to spend 80,744.